OpenStudy (el_tucan):

what is the derivative of 8 lnx. I know that the derivative of lnx is 1/x but I don't know what I do with the 8. I also know that something like lnx^2 is also 2 lnx but idk

OpenStudy (moonlitfate):

Well, since you know that \[\frac{ dy }{ dx }[\ln (x)] = \frac{ 1 }{ x }\] \[8 \ln(x) = 8 (\ln x)\] So \[\frac{ dy }{ dx } [8 \ln (x)] = 8 *\frac{ dy }{ dx }[\ln (x)]\] \[= 8 *\frac{ 1 }{ x } = \frac{ 8 }{ x }\]

OpenStudy (moonlitfate):

So \[\frac{ dy }{ dx }[8 \ln(x)] = \frac{ 8 }{ x }\]

OpenStudy (el_tucan):

sweet, thanks how do you make your work nice like that. is that math type?